OpenStudy (luigi0210):

Integration:

OpenStudy (luigi0210):

\[\LARGE \int x~(2x+5)^8~dx\]

OpenStudy (science0229):

We can use integration using substitution for this one.

OpenStudy (anonymous):

assume 2x+5=u

OpenStudy (science0229):

Right

OpenStudy (science0229):

Then, du/dx=2 So, du=2dx

OpenStudy (science0229):

\[\int\limits_{}^{}x(2x+5)^8dx=\int\limits_{}^{}xu^8*1/2*du\]

OpenStudy (science0229):

Now, from the substituion u=2x+5, we get x=(u-5)/2

OpenStudy (science0229):

\[\int\limits_{}^{}(\frac{ 1 }{ 2 })(\frac{ u-5 }{ 2 })u^8du=\int\limits_{}^{}\frac{ u^9-5u^8 }{ 4 }du\]

OpenStudy (science0229):

Can you take it from here? @Luigi0210

OpenStudy (luigi0210):

Yes, I got the rest, thanks for the explanation as well :)

OpenStudy (science0229):

Yep! Just make sure to leave your answer in terms of x, not u

sammixboo (sammixboo):

Math @-@